OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to the line 2x -3y = 12 and passes through the point (2, 6).

OpenStudy (anonymous):

The equation of the original line is y = 2x-4; this equation is in the form y = mx+c, where m is the slope of the line. As a result, the slope of the line is 2. The question requires the finding of the equation of the line that is perpendicular to the original line; as a result, we have to find the slope of the perpendicular line first. Since the slope of a line * the slope of its perpendicular line = -1; therefore, 2 * the slope of the perpendicular line = -1 the slope of the perpendicular line = -1/2. Given the perpendicular line passes through the point (-6,2), by using point slope form, which is y-y1 = m(x-x1) By substituting the point (-6,2) into the point slope form, y-2 = (-1/2)*[x-(-6)] y-2 = (-1/2)*(x+6) y-2 = (-1/2)(x)-3 y = (-1/2)(x)-1; Therefore, the slope-intercept form of the perpendicular line is y = (-1/2)(x)-1. source: http://answers.yahoo.com/question/index?qid=20081114094710AAQ5XEz

OpenStudy (zepp):

\[2x-3y=12\]\[2x-12=3y\]\[\frac{2x-12}{3}=y\]\[\frac{2x}{3}-\frac{12}{3}=y\]\[\frac{2x}{3}-4=y\]

OpenStudy (zepp):

In this equation, the slope is \[\frac{2}{3}\] The slope of the function that would be perpendicular to this function would be\[\ -\frac{3}{2}\]

OpenStudy (anonymous):

Thank you all(: